OpenStudy (anonymous):

where is tan radical 3 on the unit circle

OpenStudy (anonymous):

What is tangent?

OpenStudy (anonymous):

like sin cos tan

OpenStudy (anonymous):

I know, but what is tangent equal to? I know the answer, but I am asking you.

OpenStudy (anonymous):

i dont know:(

OpenStudy (anonymous):

Take a guess.

OpenStudy (anonymous):

just think. you know that on a 30 60 90 you have radical 3 as 1 side and 1 as 1 side. radical 3 is bigger than 1. so if rad 3 is opposite the angle and 1 is adjacent. what angle does it have to be.

OpenStudy (anonymous):

Drawing a picture might help her ^

OpenStudy (anonymous):

him or her, I don't know. >.<

OpenStudy (anonymous):



OpenStudy (anonymous):

30 degrees then?? honestly i have no idea

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

You could also go the route that I learned, \[\tan x=\frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

120 degrees??

OpenStudy (anonymous):

Ooh. Close. Try halving it. ^

OpenStudy (anonymous):

so 60 degrees??

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

We know this because:\[\tan x=\frac{ \sin x }{ \cos x }\] And at 60 degrees:\[\sin = \frac{ \sqrt{3} }{ 2 }\] and \[\cos = \frac{ 1 }{ 2 }\] \[\tan 60=\frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } }\] dividing by a fraction is equal to multiplying by it's recipricol. \[\tan 60=\frac{ \sqrt{3} }{ 2 } \times 2\] \[\tan 60=\sqrt{3}\]